Two Polaroids are placed into the unpolarized beam of a light source. The angle between
the axes of polarization of the two Polaroids is 30°. If the intensity of the incident beam is
Io, what is the intensity of the transmitted light?
1
Expert's answer
2020-11-02T09:23:36-0500

The intensity of the transmitted light is given by the Malus's law:


"I = I_0\cos^2\theta"

where "I_0" is the intensity of the incident beam and "\theta = 30\degree" is the angle between the axes of polarization of the two Polaroids. Thus, obtain:

"I = I_0\cos^2\theta = I_0\cos^230\degree = I_0 \left( \dfrac{\sqrt3}{2} \right)^2 = \dfrac{3}{4}I_0"

Answer. "I = \dfrac{3}{4}I_0".